OpenStudy (anonymous):

express the area A of a rectangle as a function of the length x if the length of the rectangle is twice its width

OpenStudy (anonymous):

since x = 2w, you can solve that for w, so you can w = x/2. to express the area as a function of length x, you will get A = x * x/2 which gives A = 1/2 x^2 !

OpenStudy (anonymous):

does that say A=x times x/2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok thanks